OpenStudy (anonymous):

Find the volume of the region bounded in back by the plane x=0 , on the front and sides by the parabolic cylinder x=1−y2 , on top by the paraboloid z=x2+y2 , and on the bottom by the xy-plane. Find the volume of the region bounded in back by the plane x=0 , on the front and sides by the parabolic cylinder x=1−y2 , on top by the paraboloid z=x2+y2 , and on the bottom by the xy-plane. @Mathematics

OpenStudy (anonymous):

do you have to integrate or just set up?

OpenStudy (anonymous):

the only thing that i am having touble with is finding out what the top bound for radius in polar coordinates thats all i need

OpenStudy (anonymous):

then i can do the rest

OpenStudy (anonymous):

I wouldn't do polar. You will have to break it up

OpenStudy (anonymous):

no you wouldnt. you never have to break things up in polar. thats only in Cartesian

OpenStudy (anonymous):

well the radius is defined as two different functions at different points. I've never heard of a way to do that in one go.

OpenStudy (anonymous):

thats if you convert it from y as a function of x to x as a function of y

OpenStudy (anonymous):

\[\int\limits_{-\pi/2}^{\pi/2}\int\limits_{0}^{?}\int\limits_{0}^{r^2}\]

OpenStudy (anonymous):

all i need is the top bound

OpenStudy (anonymous):

are you trying to use cylindrical coordinates of a non cylinder?

OpenStudy (anonymous):

should be rdrd(theta) times a height z or dz

OpenStudy (anonymous):

yeah thats what that integral is up there

OpenStudy (anonymous):

i just need to know top R

OpenStudy (anonymous):

a varying r turns it into spherical coordinates which is a different ballgame in this case. when you hit the parabaloid up top, the far radius changes

OpenStudy (anonymous):

no thats circular. The R is taken from the shadow cast on the xy field

OpenStudy (anonymous):